OpenStudy (anonymous):

A jar contains 5 red marbles, 2 blue marbles, 4 yellow marbles, and 4 green marbles. If two marbles are chosen at random, what is the probability that they will be the same color?

OpenStudy (anonymous):

Red is most likely to get the same two colors :)) btw not positive if you keep reading what the question is and the numbers of colors and marbles you will get it c:

OpenStudy (anonymous):

Red is what I would say I'm not 100% positive :)

OpenStudy (anonymous):

The probability of the marbles being the same color would be 15%. 5+2+4+4=15. Therefore, 15% is correct.

OpenStudy (anonymous):

Yep JoannThomas is right

OpenStudy (welshfella):

you need to list the favourable results that is 2 reds , 2 yellows , 2 greens or 2 blues

OpenStudy (welshfella):

P(2 reds) = 5/15 * 4/14 = 1/3 * 2/7 = 2/21

OpenStudy (welshfella):

now find the probability of getting 2 of the other colours and add the results.

OpenStudy (welshfella):

It is not 15%